OpenStudy (anonymous):

Use the squared Identities to simplify 2cos^2 (x) * cos^2 (x)

OpenStudy (anonymous):

@ganeshie8 @JackJordan

OpenStudy (anonymous):

uhmm... how simple do you want it I mean an easy one would be just 2cos^4(x) lol

OpenStudy (anonymous):

a) 1-cos(4x) / 4 b) 1+4cos(2x)+cos(4x) / 4 c) 1+cos(4x) / 4 d) 1+4cos(2x)-cos(4x) / 4

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

are you sure you haven't miswritten the second one?

OpenStudy (anonymous):

b) 1+4cos(2x)+cos(4x) / 4

OpenStudy (anonymous):

thats what i have

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@JackJordan

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

someone help!

hartnn (hartnn):

write both \(\cos^2 x = \dfrac{1-\cos 2x}{2}\)

OpenStudy (anonymous):

It must be a typo... I'm getting 1+4cos(2x)+cos(4x) + 3 / 4 sorry.

hartnn (hartnn):

***\(\cos^2 x = \dfrac{1+\cos 2x}{2}\)

OpenStudy (anonymous):

elementary mistake hartnn

hartnn (hartnn):

and then expand it.

OpenStudy (anonymous):

So my attempt leads to something similar to B but with +3 at the end. Either I'm wrong (not likely) or the question is wrong.

hartnn (hartnn):

then using the same identity you'll need to write \(\Large \cos^2 2x = \dfrac{1+\cos 4x}{2}\)

OpenStudy (anonymous):

ROOKIE MISTAKE!!!

OpenStudy (anonymous):

sorry to everyone. I am so sorry

OpenStudy (anonymous):

So I was right

hartnn (hartnn):

so you got it now?

OpenStudy (anonymous):

i must've made a huge typo

OpenStudy (anonymous):

the question is 2sin^2 (x) cos^2x

OpenStudy (anonymous):

sin. not cos. i can't believe i made that mistake.

hartnn (hartnn):

okk, then write \(\large \sin^2 x = \dfrac{1-\cos 2x}{2}\) thats the only difference

OpenStudy (anonymous):

Dang it I knew you made that mistake! Lol, it gives a difference of two squares at the top. I already did that. The answer is a. Can you see why?

hartnn (hartnn):

I'd like to see sbsbrand's attempt! :)

OpenStudy (anonymous):

yay it worked! thanks everyone that stayed to help me even tho i wrote the question wrong.......XD